### v1

# Symphony-MoE: Harmonizing Disparate Pre-trained Models into a Coherent Mixture-of-Experts

Abstract
Mixture-of-Experts (MoE) models enable scalable performance by activating large parameter sets sparsely, minimizing computational overhead. To circumvent the prohibitive cost of training MoEs from scratch, recent work employs upcycling, reusing a single pre-trained dense model by replicating its feed-forward network (FFN) layers into experts. However, this limits expert diversity, as all experts originate from a single pre-trained dense model. This paper addresses this limitation by constructing powerful MoE models using experts sourced from multiple identically-architected but disparate pre-trained models (e.g., Llama2-Chat and Code Llama). A key challenge lies in the fact that these source models occupy disparate, dissonant regions of the parameter space, making direct upcycling prone to severe performance degradation. To overcome this, we propose Symphony-MoE, a novel two-stage framework designed to harmonize these models into a single, coherent expert mixture. First, we establish this harmony in a training-free manner: we construct a shared backbone via a layer-aware fusion strategy and, crucially, alleviate parameter misalignment among experts using activation-based functional alignment. Subsequently, a single lightweight stage of router training coordinates the entire architecture. Experiments demonstrate that our method successfully integrates experts from heterogeneous sources, achieving an MoE model that significantly surpasses baselines in multi-domain tasks and out-of-distribution generalization.
1 Introduction
Figure 1: Comparison between the workflow of naive upcycling and ours.
Large language models (LLMs) have demonstrated remarkable progress in recent years, with the Mixture-of-Experts (MoE) architecture emerging as a key paradigm for efficient scaling (Shazeer et al. 2017). By dynamically routing computations to sparse subsets of parameters—known as “experts”—MoE models achieve vast total capacity while maintaining a near-constant floating-point operations (FLOPs) count per forward pass.
Despite their potential, training MoE models from scratch requires prohibitive computational resources and massive datasets. To mitigate this, academia and industry have explored upcycling techniques (Komatsuzaki et al. 2022; Zhu et al. 2024). These methods typically expand a pre-trained dense model’s feed-forward network (FFN) layers into MoE layers, initializing each expert as a replica of the original FFN. Subsequent continual training then encourages expert specialization, with studies confirming that diverse data subsets can effectively differentiate replicated experts (Sukhbaatar et al. 2024).
However, existing upcycling approaches such as “Branch-Train-Mix” (Sukhbaatar et al. 2024) share a fundamental limitation: reliance on a single pre-trained checkpoint (see Figure 1). Consequently, all experts originate from identical parameter space, regardless of subsequent specialization. This “single-origin” constraint intrinsically caps expert diversity. In practice, a large number of high-quality, identically architected models—specialized through distinct pre-training objectives, data corpora, or versions—exist, such as the dialogue-optimized Llama2-Chat (Touvron et al. 2023) and the code-centric Code Llama (Roziere et al. 2023). These models capture unique knowledge and capabilities within their domains, occupying valuable yet heterogeneous regions of the parameter space. Directly leveraging such specialized models as distinct experts offers a promising avenue for constructing more powerful and versatile MoEs, but remains infeasible under current frameworks.
Figure 2: Overview of the Symphony-MoE construction pipeline. In Stage 1, disparate source models are harmonized without training: non-FFN layers are merged into a shared backbone using techniques such as SLERP, while FFN layers are aligned through activation-based neuron permutation. In Stage 2, only the router is trained while keeping the backbone and expert parameters frozen, enabling coordination among the now-compatible experts.
Thus, this paper addresses a core research question: Can we effectively assemble MoE models by upcycling experts from disparately initialized source models? Naive upcycling encounters severe parameter space misalignment (Ainsworth, Hayase, and Srinivasa 2022). Models trained with divergent histories tend to occupy incompatible regions of both the numerical and semantic parameter space. Simply concatenating such models leads to catastrophic interference, as the misaligned experts disrupt each other’s internal representations. As a result, the router can no longer perform meaningful dispatching across functionally incoherent experts.
We propose Symphony-MoE, a framework for harmonizing a collection of disparate models into a single, powerful MoE. We achieve this harmonization through a two-stage process, where the parameter space misalignment is resolved via training-free functional alignment. In this first stage, we construct a coherent shared backbone using layer-aware fusion. We apply Spherical Linear Interpolation (SLERP) (Wortsman et al. 2022) to self-attention modules to maintain their geometric integrity, while using a selective linear strategy for token embeddings to handle vocabulary mismatches and simple averaging for the statistically-driven LayerNorm parameters. Crucially, to harmonize the experts themselves, we functionally align them to a shared representational space. This is achieved not by averaging, but by a precise reordering of neurons based on activation similarities, solved via a Linear Assignment Problem. This training-free process yields a pool of diverse yet functionally compatible experts, setting the stage for efficient coordination.
The second stage implements lightweight global coordination. We freeze the expert layers and shared backbone to train the router independently. To prevent expert collapse and ensure balanced utilization, this stage critically employs an auxiliary loss function promoting load balancing across experts.
The core contributions of this work are threefold:
(1) We investigate the problem of upcycling multiple identically architected yet independently pretrained models into a unified Mixture-of-Experts (MoE) system. We identify parameter space misalignment as a key barrier to this process..
(2) We introduce a lightweight framework that addresses expert incompatibility without gradient-based training, combining layer-aware fusion and activation-driven functional alignment to preserve the specialized knowledge embedded in each source model.
(3) Through comprehensive experiments on in-distribution and out-of-distribution datasets, we show that our method achieves strong performance and generalizes well to out-of-distribution settings, demonstrating the feasibility of upcycling disparate pre-trained models into unified expert systems.
2 Related Works
2.1 Mixture-of-Experts Models
The Mixture-of-Experts (MoE) paradigm is originally proposed as a “divide and conquer” modular learning system (Jacobs et al. 1991). Its modern renaissance was triggered by the Sparsely-Gated MoE (Shazeer et al. 2017), which introduced sparse, top-k routing to enable conditional computation at an unprecedented scale. This decouples the model’s parameter count from its computational cost.
Subsequent advances, most notably the Switch Transformer (Fedus, Zoph, and Shazeer 2022), simplified the routing mechanism to top-1 selection, improving training stability and reducing communication overhead. This streamlined approach established sparsity as a core strategy for scaling, as adopted in models like GLaM (Du et al. 2022) and Mixtral (Jiang et al. 2024). However, training these massive MoE models from scratch remains prohibitively expensive, which motivates more resource-efficient construction methods.
2.2 MoE Construction via Upcycling
To mitigate the immense cost of training MoEs from scratch, the upcycling paradigm has emerged as a practical alternative. The vanilla approach, or Sparse Upcycling (Komatsuzaki et al. 2022; He et al. 2024), involves converting a pre-trained dense model into an MoE by replicating its feed-forward network (FFN) layers to form multiple, initially identical experts. While this provides a strong initialization, it suffers from a knowledge bottleneck and a lack of initial expert diversity, as all experts originate from a single source.
To address this, Drop-Upcycling (Nakamura et al. 2025) performs partial re-initialization to enhance expert diversity. “Branch-Train-Mix” strategy (Sukhbaatar et al. 2024; Horoi et al. 2025; Kang et al. 2024; Li et al. 2024a) constructs an MoE from multiple, pre-specialized dense models to inject diverse knowledge from the outset. For example, BAM (Zhang et al. 2024) first fine-tunes copies of a base model on different domains and then uses their specialized FFN and attention layers as the experts in the final MoE. However, it introduces a new critical challenge: how to effectively fuse the knowledge from these disparate trained specialists, whose parameters reside in different regions of the loss landscape. Current methods like BAM and Self-MoE (Kang et al. 2024) rely on rudimentary fusion techniques, such as simple averaging for non-expert layers, which are not guaranteed to be optimal.
2.3 Model Merging
The challenge of combining disparate pre-trained models is the central focus of the model merging field (Tao et al. 2024). To merge truly “stranger” models, advanced techniques explicitly resolve misalignment. Model Soups (Wortsman et al. 2022) averages the weights of models fine-tuned from the same checkpoint, assuming they remain in the same well-connected loss basin. Methods like PLeaS (Ito, Yamada, and Kumagai 2024) and Git Re-Basin (Ainsworth, Hayase, and Srinivasa 2022) find an optimal permutation matrix to align neurons layer-wise before averaging. TIES-Merging (Yadav et al. 2023) resolves interference between task vectors by trimming redundant updates and electing a dominant sign for parameter changes. These methods, however, are general-purpose algorithms designed to produce a single, dense consensus model. They are not tailored for the specific architectural needs of MoE construction, such as preserving expert diversity or informing the initialization of the router.
Our work bridges this gap. We integrate permutation-based alignment directly into the MoE construction pipeline, repurposing these techniques for a new architectural goal. We use alignment to make disparate experts functionally compatible. This ensures they can operate within a shared coordinate system while preserving their individual specializations.
3 Methods
We introduce a two-stage framework to effectively assemble a MoE model from disparate, pre-trained source models. The first stage operates entirely without training, resolving the critical parameter space misalignment problem through training-free functional alignment. The second stage then trains a router that coordinates the now-compatible experts, unlocking their collective capabilities.
3.1 Preliminaries and Problem Formulation
We first establish the requisite notation for Transformer and MoE architectures and then formally define the problem.
A standard Transformer model is composed of stacked layers, each typically containing a multi-head self-attention mechanism followed by a position-wise feed-forward network (FFN). In an MoE architecture, the dense FFN layer is replaced by a set of $N$ independent “expert” networks $\{E_{1},E_{2},\ldots,E_{N}\}$, each with the same architecture as the original FFN. A trainable router network, $R$, directs the input tokens to a sparse subset of these experts. For each input token $x$, the router computes gating values $g(x)$ that determine how the outputs of the experts are combined:
$M_{MoE}(x)=\sum_{i=1}^{N}g_{i}(x)\cdot E_{i}(x)$ (1)
The gating values are typically produced by a linear layer followed by a softmax function, $g(x)=Softmax(xW_{g})$, where $W_{g}$ is a trainable routing matrix. For sparsity, often only the top-k experts with the highest gating values are activated for any given token.
Our work addresses the novel challenge of constructing an MoE model from multiple, disparate pre-trained source models. Formally, we are given a set of $N$ source models, $\{M_{1},M_{2},\ldots,M_{N}\}$. These models share an identical architecture but possess distinct sets of parameters $\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$ as a result of being trained on different datasets $\{D_{1},D_{2},\ldots,D_{N}\}$ , for different tasks, or with different optimization objectives. The goal is to construct a single, powerful MoE model, $M_{MoE}$, by leveraging the FFN layers of the source models as its experts. The fundamental barrier to this goal is the severe parameter space misalignment between the models, where the parameters $\theta_{i}$ and $\theta_{j}$ (for $i\neq j$) occupy incompatible numerical and semantic spaces.
3.2 Stage 1: Training-Free Functional Alignment
Our first stage assembles a cohesive MoE model from diverse parent models entirely without training. This process involves two key steps: fusing the shared architectural backbone and functionally aligning the FFN layers, which will serve as the experts.
Shared Backbone Construction. We first construct the shared, non-expert layers. Our approach to this fusion is explicitly layer-aware. Rather than applying a single, uniform merging technique, we tailor the fusion strategy to the specific architectural function of each component. Specifically:
$\cdot$ For the token embedding layers, we follow the selective linear strategy implemented in MergeKit (Goddard et al. 2024). For tokens shared across models, the algorithm retrieves their corresponding embedding vectors from each model and performs standard linear averaging. For tokens that are unique to a specific model—i.e., not present in the vocabularies of others—the original embedding is preserved without modification.
$\cdot$ For self-attention modules, which are critical for contextual understanding, we apply Spherical Linear Interpolation (SLERP) (Wortsman et al. 2022). This method better preserves the geometric integrity of the weight space and mitigates the functional degradation often caused by naive linear averaging. We apply it to all weight matrices (Q, K, V, O projections).
$\cdot$ In contrast, the primary statistical function of LayerNorm parameters makes a simple linear average a sufficient and stable choice  (Jin et al. 2022).
Pairwise Inter-Expert Alignment. We designate one model’s FFNs as the “anchor” and then permute the neurons of the other models’ FFNs to match the anchor’s functional behavior based on activation patterns.
For each non-anchor model $M_{i}$ and each layer $l$, the alignment process is as follows:
1. Activation Collection: We construct a small and diverse calibration set $D_{\mathrm{cal}}$ by sampling 5% from each instruction fine-tuning dataset $D_{1},D_{2},\ldots,D_{N}$} used in dense model training. Each instruction in $D_{\mathrm{cal}}$ is passed through both the anchor model $M_{1}$ and a target model $M_{i}$ to extract their FFN activation matrices at layer $l$, denoted as $A_{1}^{(l)}$ and $A_{i}^{(l)}$, respectively.
2. Permutation Matching: We find the optimal permutation matrix $P_{i}^{(l)}$ that aligns the neurons of $M_{i}$ to $M_{1}$ by solving the linear assignment problem:
[…]
where $\mathcal{P}$ is the set of permutation matrices and $||\cdot||_{F}$ is the Frobenius norm. This problem can be solved efficiently with the Hungarian algorithm (Mills-Tettey, Stentz, and Dias 2007).
3. Weight Remapping: We apply the computed permutation $P_{i}^{(l)}$ to the FFN weights (Nasery et al. 2025). This involves permuting the output dimension of the first linear layer’s weights $W_{up}$ and the input dimension of the second’s $W_{down}$ to maintain functional equivalence:
$W^{\prime}_{up,i}=W_{up,i}^{(l)}P_{i}^{(l)},\quad W^{\prime}_{down,i}=(P_{i}^{(l)})^{T}W_{down,i}^{(l)}$ (3)
This procedure results in $N$ functionally aligned experts for each layer, all operating in a consistent parameter space.
We analyzed the computational cost to verify the scalability. The analysis can be found in the Appendix.
3.3 Stage 2: Activating the MoE via Lightweight Router Training
In Stage 2, we integrate the constructed shared backbone with the expert models and introduce a randomly initialized router, implemented as a simple linear layer. The expert layers and shared backbone—which encode knowledge from the source models—are frozen, and only the router is trained. The calibration dataset $D_{\mathrm{cal}}$, introduced in Stage 1, is reused for this purpose, enabling the router to learn how to coordinate the experts.
We adopt a top-2 routing mechanism and train the router using a composite loss function:
$\mathcal{L}_{total}=\mathcal{L}_{lm}+\lambda\cdot\mathcal{L}_{bal}$ (4)
where $\mathcal{L}_{lm}$ is the standard causal language modeling objective, and $\mathcal{L}_{bal}$ is a load-balancing regularization term. The coefficient $\lambda$ is fixed at 0.01 in all experiments. $\mathcal{L}_{bal}$ encourages uniform expert utilization, which mitigates expert collapse and allows the model to fully leverage the diversity of its specialists.
Notably, the router is not provided with explicit supervision indicating which expert to select for each token, as there is no ground-truth mapping between tokens and experts. Instead, it learns expert assignment implicitly through training on instruction–response pairs in $D_{\mathrm{cal}}$.
4 Experiments
In this section, we aim to address the following key research questions: (1) How does Symphony-MoE perform at different model scales or types, compared to baseline models on both in-distribution and out-of-distribution (OOD) datasets? (2) Does our proposed alignment method effectively mitigate parameter space misalignment across experts? (3) How does the choice of anchor model influence the final performance of Symphony-MoE? (4) What is the impact of varying the number of experts on model performance? (5) Ablation Analysis: How sensitive is the system to variations in alignment and merging strategies?
4.1 Experimental Setup
4.1.1 Build Disparately Initialized Dense Models
ID OOD
MMLU GSM8K BBH HumanEval TruthfulQA Avg. MedCQA
Dense (0.5B)
General ($M_{1}$, ANC) 44.27±0.40 18.87±1.07 27.46±0.50 24.40±3.36 27.29±1.56 28.46 26.20±0.19
Math ($M_{2}$) 42.07±0.40 19.33±1.08 22.72±0.46 20.73±3.17 28.40±1.58 26.65 26.23±0.19
Code ($M_{3}$) 30.96±0.38 20.48±0.41 25.33±0.49 28.71±3.27 25.81±1.55 26.26 25.93±0.19
Science ($M_{4}$) 39.25±0.41 21.30±1.12 22.98±0.48 18.29±3.02 30.01±1.50 26.37 25.56±0.19
MoE (0.5B $\times$ 4)
BTX 37.28±0.61 18.95±0.42 25.57±0.48 26.59±3.15 23.90±1.19 26.46 24.02±0.19
BAM 42.76±0.70 19.15±0.68 27.68±0.58 26.77±3.07 26.54±1.61 28.58 26.79±0.19
Drop 44.08±0.40 22.85±0.39 26.51±0.48 25.52±3.19 30.81±1.55 29.95 27.95±0.20
Symphony (Ours) 45.10±0.40 24.57±1.11 29.64±0.48 28.02±3.12 31.54±1.50 31.77 29.07±0.19
Table 1: Performance comparison of dense (Qwen2/2.5 0.5B) and MoE models (upcycled from dense models) on in-distribution (ID) and out-of-distribution (OOD) data. The best and second-best results are bolded and underlined. “ANC” stands for anchor model.
ID OOD
MMLU GSM8K BBH HumanEval TruthfulQA Avg. MedCQA
Dense (1.5B)
General ($M_{1}$, ANC) 59.57±0.39 33.72±0.39 44.39±0.55 37.19±3.79 29.87±1.60 40.95 30.39±0.20
Math ($M_{2}$) 56.73±0.39 34.11±0.50 43.97±0.54 37.20±3.79 31.33±1.62 40.67 30.36±0.20
Code ($M_{3}$) 43.74±0.41 35.86±1.33 36.17±0.55 43.41±3.81 26.19±1.51 37.07 28.72±0.19
Science ($M_{4}$) 54.60±0.40 32.45±1.30 36.05±0.53 35.37±3.74 30.78±1.57 37.85 29.72±0.19
MoE (1.5B $\times$ 4)
BTX 45.12±0.88 30.35±0.45 40.11±0.58 29.08±3.15 25.02±1.15 33.94 26.92±0.18
BAM 50.77±0.75 36.99±0.61 45.01±0.59 37.84±2.89 27.14±1.89 39.55 28.97±0.18
Drop 57.14±0.81 34.92±1.12 44.83 ±0.43 36.72±1.89 31.88±1.49 41.10 32.90±0.19
Symphony (Ours) 58.91±0.31 39.12±1.09 46.97±0.36 42.39±3.69 32.95±2.31 44.07 35.26±0.22
Table 2: Performance comparison of dense (Qwen2/2.5 1.5B) and MoE models (upcycled from dense models) on in-distribution (ID) and out-of-distribution (OOD) data. The best and second-best results are bolded and underlined. “ANC” stands for anchor model.
To ensure a rigorous evaluation of our proposed merging method, we first constructed a set of four disparately initialized dense models at scale of 0.5/1.5B. This experimental suite was designed to reflect real-world disparity by systematically incorporating variations across three dimensions: model versions, pre-training data domains, and downstream instruction-tuning tasks.
Our framework is built upon three distinct foundation checkpoints, each contributing a different source of disparity. The first is Qwen2.5-Base, a general-purpose model trained on broad-domain natural language data, with a focus on tasks involving understanding and generation. To introduce task- and data-level disparity, we include Qwen2.5-Coder (Hui et al. 2024), which is primarily trained on code-related corpora and optimized for tasks such as code completion and code–natural language interconversion. Finally, to incorporate version-level disparity, we use Qwen2-Base (Team 2024), an architecturally compatible predecessor of Qwen2.5-Base with a completely different pre-training history.
Based on the three foundational models, we derived four specialist experts through instruction fine-tuning. The generalist expert ($M_{1}$) and the mathematics expert ($M_{2}$) were obtained by fine-tuning Qwen2.5-Base on the Alpaca (Maeng, Colin, and Lucia 2017) and MetaMathQA (Yu et al. 2024) datasets, respectively. The code expert ($M_{3}$) was derived from our code-specialized foundation using the CodeAlpaca (Li et al. 2024b) dataset. The science expert ($M_{4}$) was created by fine-tuning Qwen2-Base on the SciQAG (Wan et al. 2024) dataset.
Further training details are provided in the Appendix.
4.1.2 Evaluation Datasets and Metrics
We evaluate our method on five in-distribution datasets: MMLU (General)  (Hendrycks et al. 2021), GSM8K (Math) (Cobbe et al. 2021), BBH (General, Reasoning)  (Suzgun et al. 2022), HumanEval (Code) (Chen et al. 2021), and TruthfulQA (Science) (Lin, Hilton, and Evans 2022). We also include an out-of-distribution benchmark, MedCQA (Medicine) (Shoham and Rappoport 2024).
For evaluation, we follow the standard metrics used in each benchmark: MMLU and MedCQA are evaluated using zero-shot accuracy, GSM8K and BBH use eight-shot accuracy, HumanEval is measured with pass@1, and TruthfulQA adopts the MC1 metric.
4.1.3 Baselines
We compare our upcycling approach against two categories: the dense models $M_{1}$, $M_{2}$, $M_{3}$, $M_{4}$ that we upcycle, and recent upcycling methods (BTX (Sukhbaatar et al. 2024), BAM (Zhang et al. 2024), Drop (Nakamura et al. 2025)). To ensure fairness, the baseline method uses the same data as ours for fine-tuning the dense model. We reproduced the results according to the original paper, and the reproduction details are in the appendix.
4.2 Implementation Details
Figure 3: Quantitative analysis of inter-expert functional specialization using Centered Kernel Alignment (CKA). Lower CKA scores indicate greater functional specialization. High CKA scores between experts reflect parameter space misalignment, caused by failing to align neurons functionally during merging. This misalignment leads to representational collapse, erasing the distinct capabilities of individual experts.
Our experiments were conducted on 8 NVIDIA V100 GPUs. The dense model was implemented using LLaMA-Factory (Zheng et al. 2024), and trained for for 2 epoches. Detailed training configurations are provided in the Appendix.
In Stage 1: Alignment, we randomly sampled from each training dataset to construct the calibration set $D_{\mathrm{cal}}$, resulting in a total of 10.404 million tokens. To ensure consistent alignment and integration across expert models, we fixed a general-purpose expert model $M_{1}$ as the anchor model. To ensure that the knowledge of each expert is equally integrated and to avoid introducing any bias, the weights for merging the 4 dense models are set to 0.25.
In Stage 2: Router Training, we set the cutoff length to 2048 and train the model for 6 epochs with a batch size of 2. The learning rate is fixed at 5e-5, and optimization is performed using AdamW (Zhuang et al. 2022) with a maximum gradient norm of 1.0.
In the evaluation phase, we used lm-evaluation-harness-v0.4 (Gao et al. 2024) to obtain scores for datasets such as MMLU and GSM8K.
4.3 Main Results
Overall Performance. As summarized in Table 1 and 2, Symphony-MoE achieves the highest average score across all evaluation datasets, substantially outperforming all MoE baselines, including BTX, BAM, and Drop. In contrast, individual dense models demonstrate strong domain specialization but suffer severe performance degradation on out-of-domain tasks, indicating limited generalization.
In-Domain Expertise Preservation. Importantly, Symphony-MoE is not a naive average of its experts. On in-domain tasks, it effectively retains the specialized capabilities of its constituent models. For example, in the HumanEval, although the dedicated code expert $M_{3}$ achieves the top score, Symphony-MoE performs comparably, trailing by only 1–2 percentage points.
Out-of-Distribution Generalization. Beyond domain retention, Symphony-MoE demonstrates superior generalization ability, outperforming all baselines on MedCQA. This suggests that our activation-based alignment strategy enables the model to internalize more transferable and fundamental reasoning patterns, rather than merely interpolating between expert skills.
Robustness Across Scales. We further validate these findings across two model scales—4 experts × 0.5B and 4 experts × 1.5B. In both settings, Symphony-MoE consistently exhibits strong average performance, competitive in-domain expertise, and leading out-of-distribution generalization, confirming the robustness of our approach.
Robustness Across Types. We further validated the effectiveness of Symphony-MoE using the Llama model as the source architecture. As shown in the Appendix, our method yields consistent improvements, demonstrating its robustness across different backbone model types.
4.4 Analysis of Inter-Expert Functional Specialization
To quantitatively assess the effectiveness of our functional alignment method, we compute the pairwise functional specialization between expert layers using Centered Kernel Alignment (CKA) (Cortes, Mohri, and Rostamizadeh 2012), a widely adopted technique for comparing neural network representations. A lower CKA score indicates greater functional specialization. We evaluate three scenarios: (a) the original, unmerged expert models as a reference baseline; (b) a naive merging approach without alignment; and (c) our proposed method employing activation-based functional alignment.
As shown in Figure 3, the naive merging strategy leads to a severe representational collapse. Experts trained independently develop distinct internal coordinate systems; merging them without alignment forces the model to conflate functionally unrelated neurons. This results in inflated inter-expert CKA scores (e.g., 0.65–0.75), signaling that unique functional fingerprints of the experts have been blurred into a redundant subspace. In such a scenario, the router lacks the discriminative basis to perform meaningful, specialized dispatching.
In contrast, our method explicitly resolves this misalignment by permuting neurons to form a functionally coherent coordinate system prior to merging. Figure 3(c) demonstrates that this alignment dramatically restores inter-expert specialization, with CKA scores returning to near-optimal levels comparable to those of the original experts (Figure 3(a)). This provides empirical evidence that our alignment strategy effectively mitigates parameter space misalignment.
4.5 Analysis of Anchor Model Selection
Figure 4: Performance analysis of Symphony-MoE. (a) Average in-distribution (ID) and out-of-distribution (OOD) scores under different anchor model choices. (b) Impact of increasing the number of experts from 1 to 4 on ID and OOD performance.
[…]
As shown in the Figure 4(a), using the general-purpose expert model $M_{1}$ as the anchor yields the highest average scores on both ID and OOD tasks. This suggests that a balanced and non-specialized functional space plays a critical role in constructing an effective “omnipotent” MoE. It facilitates collaboration among experts without constraining them to overly narrow subspaces.
4.6 Analysis of the Number of Experts
[…]
4.7 Ablation Analysis
Model / Configuration ID Avg. OOD
(+) Full Method 44.07 35.26
(-) No Functional Alignment 33.94 26.92
(-) Naive Attention Merge 37.28 31.77
(-) Naive Embedding Merge 40.19 31.95
(-) Biased Calibration Data 39.70 29.44
Table 3: Ablation analysis of key components in Symphony-MoE (1.5B x 4).
We assessed the impact of key components on the performance of the Symphony-MoE model (1.5B × 4 configuration) through an ablation study, where each core module was individually removed or replaced. The results are presented in Table 3. Specifically, Naive Attention Merge replaces the SLERP fusion strategy with simple linear averaging when constructing the shared backbone for self-attention layers in Stage 1. This setting quantifies the contribution of SLERP to preserving the functional integrity of attention mechanisms. Naive Embedding Merge replaces MergeKit with a naive linear averaging of overlapping token embeddings during the construction of the shared embedding layer. Biased Calibration Data alters the sampling strategy of the calibration dataset $D_{\mathrm{cal}}$ in the alignment stage, using samples exclusively from the training set of the general expert $M_{1}$ rather than from all expert domains.
The results highlight the critical role of functional alignment: removing it causes catastrophic performance collapse, confirming that resolving parameter space inconsistencies is fundamental to the framework. Replacing SLERP with linear averaging leads to substantial degradation, supporting our hypothesis that the geometry of attention weights is essential for preserving functionality. While simplifying the embedding fusion yields only a moderate drop, it nonetheless indicates the value of layer-aware merging in the shared backbone. Notably, using biased calibration data results in the most severe performance decline after removing alignment, underscoring the importance of domain diversity in constructing $D_{\mathrm{cal}}$ for effective inter-expert coordination.
5 Conclusion
This paper presents Symphony-MoE, a framework for building powerful Mixture-of-Experts (MoE) models by upcycling experts from multiple, diverse pre-trained sources. We tackle the core challenge of parameter space misalignment through a two-stage process: training-free functional alignment and lightweight router training. Experiments show that Symphony-MoE consistently outperforms baselines on multi-domain tasks and generalizes well to out-of-distribution data. This approach offers a scalable path to leverage the collective knowledge embedded in existing specialized models—beyond the limitations of single-model upcycling.
Limitations and future work. Our current framework requires that all source models share an identical architecture. This constraint limits the pool of potential experts that can be integrated. Future work could focus on developing more advanced alignment techniques capable of harmonizing models with minor architectural dissimilarities, such as different activation functions or normalization strategies.
The quality of the functional alignment is contingent on the diversity and representativeness of the calibration dataset. While we proposed a simple and effective sampling strategy, optimizing the curation for a highly diverse set of experts is a non-trivial challenge.